Suppose total benefits and total costs are given by B(Y) = 600Y − 12Y 2 and C(Y) = 20Y 2. What is the maximum level of total benefits?

a. 7,500

b. 1,600

c. 2,812.5

d. None of the statements associated with this question are correct.
Answer #1

Ans) the correct option is c. 2,812.5

MB =dB/dY = 600 - 24Y

MC = dC/dY = 40Y

At Profit maximization, MB =MC

600 - 24Y= 40Y implies Y = 9.375

Putting the value of Y = 9.375

B - C = 600Y -12Y2 - 20Y2 = 600Y - 32Y2

600Y - 32Y2 = 2812.5